# HandSSCA: 3D Hand Mesh Reconstruction with State Space Channel Attention from RGB images

### v2

@@ 3.4 Regressor
In Regressor, we input the hand attention features obtained from SSMA. First, a 2D heat map of each joint $H$ is output through a simple Conv cascade. Then, a cascade of hand attention features and 2D heatmap $HM$ is input into four residual blocks [34]. Finally, the outputs of the residual blocks are vectorized into 2048-dimensional vectors and passed to the fully-connected layer for predicting the pose parameter $\theta\in\mathbb{R}^{48}$ and shape parameter $\mathfrak{\beta}\in\mathbb{R}^{10}$ of the MANO. At the same time, the joint regression matrices are multiplied by the original 3D mesh $\overline{\bar{T}}$ , and the forward kinematics are applied to obtain the final 3D hand joint coordinates $J\in\mathbb{R}^{21\times 3}$ and the 3D hand mesh $V\in\mathbb{R}^{778\times 3}$.
To train our proposed HandSSCA hand pose estimation model, we constrain the model by calculating L2 loss on MANO parameters, 2D hand keypoints, 3D hand keypoints and vertex coordinates.
@@ 4.1 Implementation details
All implementations were done in PyTorch. We used the Adam optimizer kingma2017adam, and each model was trained by annealing every 10th time starting from an initial learning rate of 10-4 learning rate to train with a batch size of 32. Both our training and testing were performed on a server with a CPU15 vCPU AMD and an RTX A5000 (24GB). All other details will be provided in our code.
@@ 4.2 Datasets and evaluation metrics
For FreiHand, we report PA-MPJPE and PA-MPVPE in mm as well as F-scores. For HO3D we report PA-MPJPE/MPVPE, AUC and F-scores returned from the official evaluation server. For DexYCB we report MPJPE and PA-MPJPE in mm.
@@ 4.3 Comparisons with the State-of-the-art Methods
We compare our proposed method with several recent state-of-the-art methods on three datasets, FreiHand, DEXYCB, and HO3D, as shown in Tables 1-3. As can be seen from Table 1, our method maintains state-of-the-art performance in the F@15 metric on a smaller number of datasets, although it is slightly worse than the other methods in the PA-MPJPE metric cho2022crossattention; 9710830. Moreover, its number of parameters shrinks by a factor of nearly 5 compared to models with higher accuracy cho2022crossattention. On the HO3D dataset, our method reduces the number of parameters by about 40% compared to recent methods Hampali2021KeypointTS. Meanwhile, when the number of parameters is similar to the method of Liu et al. liu2021semisupervised, all other metrics are greatly improved, and only the F-score metric is slightly reduced. To verify the effectiveness of our proposed method in the case of large-scale occlusion, the metrics on the DEXYCB dataset are shown in Table 2, and all the metrics outperform the recent state-of-the-art models Wang_2023. In conclusion, the experimental results on all three datasets show that our proposed method can still be effective in estimating hand postures when a large number of occlusions are included. Moreover, the state-of-the-art performance can be maintained while reducing the number of parameters, as shown in Figure.4.
@@ 4.4 Ablation Study
In order to validate the effectiveness of the HandSSCA method, we further analyzed the different attention modules. The results of the ablation experiments on different datasets are shown in Tables 4-6. On the FreiHand and DEXYCB datasets, we designed three experiments, Baseline, Adding Self-Attention Module, and Adding SSCA Module, as a way to validate that our proposed SSCA is able to expand the effective receptive field while enhancing the hand features. On the HO3D dataset, we designed four experiments, which are Baseline, adding the self-attention module, adding the traditional state space attention module (SSA) and adding the state space channel attention module (SSCA). In this way, we demonstrate that the state space model with the addition of channel information can enhance the ability to extract hand features in the occluded state while having a wider effective receptive field.
From Tables 4 and 5, we can see that in the presence of occlusion on the hand, the global feature information obtained from the self-attention mechanism, although it can enhance the hand features to a certain extent, will introduce more noise information. In Figure.5, we visualize the attention distribution map, and it can be clearly seen that the feature map tends to characterize the global information after adding the attention mechanism, but still does not extract the hand features well for the occluded regions and hand interaction edges. The SSA module, on the other hand, aggregates more hand information and performs better at the interaction edges, but it will characterize the occluded part as the hand, which introduces interference information to a certain extent. The SSCA module expands the range of the receptive field through the spatial plane selection mechanism, and then directionally strengthens the hand region features through the selection of the channel planes. Furthermore, by benefiting from linear directional scanning, the occluded areas are characterized using adjacent hand information, which reduces the introduction of interference through this global mapping capability.
@@
Figure 4: Qualitative results for the three datasets. From left to right are the input image containing severe occlusions and angular distortions, the attention mapping map obtained by SSCA, the heat map of keypoints in Regressor, and the hand mesh.
Figure 5: Attention characterization map. It is worth noting that the SSCA module is able to obtain sufficient hand information in the hand interaction edge region and the occlusion region. And the method expands the range of the receptive field so that it can capture the global feature mapping. (Orange is the occluded region and red is the hand interaction edge region)
@@ 5 Conclusion
In order to effectively mitigate the challenges of the hand pose estimation task under occlusion conditions, we introduce state-space modeling into the field of hand pose estimation for the first time and propose the HandSSCA method. The method expands the effective receptive field and acquires hand features under occlusion by parallel cross-scanning in the spatial plane and the channel plane. Among them, the spatial plane is able to acquire feature associations in different directions for each pixel, and the channel plane activates the target region features through state selection. This novel attention architecture is able to acquire sufficient global information and expand the effective sensory field while keeping the number of parameters small. Extensive experimental results show that this method is able to accurately predict hand postures from heavily occluded situations. Meanwhile, we expect this spatial and channel-parallel state-selective attention architecture to make progress in different downstream tasks.

### v3

@@ 3.4 Regressor
In Regressor, we input the hand attention features obtained from S3C into it. First, a 2D heat map of each joint $H$ is output through a simple Conv cascade. Then, a cascade of hand attention features and 2D heatmap $HM$ is input into four residual blocks he2015deep. Finally, the outputs of the residual blocks are vectorized into 2048-dimensional vectors and passed to the fully-connected layer for predicting the pose parameter $\theta\in\mathbb{R}^{48}$ and shape parameter $\mathfrak{\beta}\in\mathbb{R}^{10}$ of the MANO. At the same time, the joint regression matrices are multiplied by the original 3D mesh $\overline{\bar{T}}$ , and the forward kinematics are applied to obtain the final 3D hand joint coordinates $J\in\mathbb{R}^{21\times 3}$ and the 3D hand mesh $V\in\mathbb{R}^{778\times 3}$.
To train our proposed HandS3C hand mesh reconstruction model, we constrain the model by calculating L2 loss on MANO parameters, 2D hand keypoints, 3D hand keypoints and vertex coordinates.
@@ 4.1 Implementation Details
All implementations were performed with PyTorch. We used the Adam optimizer kingma2017adam, and each model was trained by annealing every 10th time starting from an initial learning rate of 10-4 to train the model with batch size of 32. Both our training and testing were performed on a server with an GPU RTX A5000 (24GB). All other details will be provided in our code.
Figure 4: Qualitative results for FREIHAND and HO3D. From left to right are the input image containing severe occlusions and angular distortions, the attention mapping map obtained by S3C, the heat map of keypoints in Regressor, and the hand mesh.
@@ 4.2 Datasets and Evaluation Metrics
For FreiHand, we report PA-MPJPE and PA-MPVPE in $mm$ as well as F-scores. For HO3D, we report PA-MPJPE/MPVPE, AUC and F-scores returned from the official evaluation server. For DexYCB, we report MPJPE and PA-MPJPE in $mm$.
@@ 4.3 Comparisons with the State-of-the-art Methods
We compare our proposed method with several recent state-of-the-art methods on three datasets as shown in Tables 1-3. As can be seen from Table 1, our method maintains state-of-the-art performance in the F@15 metric on a smaller number of datasets, although it is slightly worse than the other methods in the PA-MPJPE metric cho2022crossattention; 9710830. Moreover, its number of parameters shrinks by a factor of nearly 5 compared to models with higher accuracy cho2022crossattention. On the HO3D dataset, our method reduces the number of parameters by about 40% compared to recent methods Hampali2021KeypointTS. Meanwhile, when the number of parameters is similar to the method of Liu et al. liu2021semisupervised, all other metrics are greatly improved, and only the F-score metric is slightly reduced. To verify the effectiveness of our proposed method in the case of large-scale occlusion, the metrics on the DEXYCB dataset are shown in Table 2, and all the metrics outperform the recent state-of-the-art models Wang_2023; 10204388; Feng20243DHR. In conclusion, the experimental results on all three datasets show that our proposed method can still be effective in estimating hand postures when a large number of occlusions are included. Moreover, the state-of-the-art performance can be maintained while reducing the number of parameters, as shown in Fig.4 - 5.
Figure 5: Qualitative results for DEXYCB. From left to right are the input image containing severe occlusions and angular distortions, the attention mapping map obtained by S3C, the heat map of keypoints in Regressor, the hand mesh and HandOCC park2022handoccnet hand mesh results.
Figure 6: Attention characterization map. It is worth noting that the S3C module is able to obtain sufficient hand information in the hand interaction edge region and the occlusion region. And the method expands the range of the receptive field so that it can capture the global feature mapping. (Orange is the occluded region and red is the hand interaction edge region)
@@ 4.4 Ablation Study
In order to validate the effectiveness of the HandS3C method, we further analyze the different attention modules. The results of the ablation experiments on different datasets are shown in Tables 4-6. On the FreiHand and DEXYCB datasets, we perform three experiments, i.e., Baseline, Adding Self-Attention Module, and Adding S3C Module, as a way to validate that our proposed S3C is able to expand the effective receptive field while enhancing the hand features. On the HO3D dataset, we conducted five experiments: baseline, adding a self-attention module, adding a traditional state-space attention module (SSA), adding a serial state-space spatial-channel attention module (serial-S3C), and adding a parallel state-space spatial-channel attention module (S3C). With this approach, we demonstrate that a state-space model with added channel information enhances the ability to extract hand features under occlusion, with a wider effective receptive field, and the effectiveness of the parallel structure.
From Tables 4 and 5, we can see that in the presence of occlusion on the hand, the global feature information obtained from the self-attention mechanism, although it can enhance the hand features to a certain extent, will introduce more noise information. In Fig. 6, we visualize the attention distribution map, and it can be clearly seen that the feature map tends to characterize the global information after adding the attention mechanism, but still does not extract the hand features well for the occluded regions and hand interaction edges. The SSA module, on the other hand, aggregates more hand information and performs better at the interaction edges, but it will characterize the occluded part as the hand, which introduces interference information to a certain extent. The S3C module expands the range of the receptive field through the spatial plane selection mechanism, and then directionally strengthens the hand region features through the selection of the channel planes. Furthermore, by benefiting from linear directional scanning, the occluded areas are characterized using adjacent hand information, which reduces the introduction of interference through this global mapping capability.
@@
For the choice of serial and parallel structures, success was achieved in image restoration through serial structures in the study of Deng et al. Deng2024CUMambaSS. However, since SSM resembles the linear computation of recurrent neural networks (RNN). Serial scanning in space and channel add extra computation time. Meanwhile, compared to the parallel structure, the serial structure does not balance the contributions of spatial and channel dimensions well, as shown in the ablation experiments in Table 6.
@@ 5 Conclusion
In order to effectively mitigate the challenges of the hand mesh reconstruction task under occlusion conditions, we introduced the state space model into the task of hand mesh reconstruction for the first time and propose the HandS3C method. The method expanded the effective receptive field and acquired hand features under occlusion by parallel cross-scanning in the spatial and channel planes. Among them, the spatial plane is able to acquire feature associations in different directions for each pixel, and the channel plane activates the target region features through state selection. This novel attention architecture is able to acquire sufficient global information and expand the effective sensory field while keeping the number of parameters small. Extensive experimental results show that this method is able to accurately predict hand postures from heavily occluded situations. Meanwhile, we expect this spatial and channel-wise state-space attention architecture to make progress in different downstream tasks.
[…]